OpenStudy (anonymous):

Find the equation of the vertical asymptote of the function: f(x)=4/(x+5)

OpenStudy (anonymous):

first what is an asymptote?

OpenStudy (anonymous):

oh another. guess what the domain is? all real numbers except x = -5 right?

OpenStudy (anonymous):

and that is the answer to your question exactly. the vertical asymptote is \[x=-5\]

OpenStudy (anonymous):

meaning [-5,infinity) (-infinity,-5]

OpenStudy (anonymous):

hold on i will get a picture and you can see it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

never write that you teacher will think you are sleeping. write \[(-\infty,-5)\cup (-5,\infty)\] smaller on the left. and round brackets because you are excluding -5

OpenStudy (anonymous):

you can see that at -5 the function either goes up to positive infinity down to negative infinity. that is the asymptote

OpenStudy (anonymous):

a vertical asymptote of f(x)/g(x) is when g(x) = 0

OpenStudy (anonymous):

you can see it even clearer here http://www.wolframalpha.com/input/?i=y%3D4%2F%28x%2B5%29++++x+from+-20+to+10

OpenStudy (anonymous):

that vertical line is not actually part of the graph, but that is what the machine makes when it connects the dots. function approaches that vertical line. but remember, asking for the vertical asymptote is usually exactly the same as what you did before. set the denominator equal to zero and solve

OpenStudy (anonymous):

ok, i really dont like this part of math. its hard...lol.

OpenStudy (anonymous):

would i be answering this the same way that i would the other questions that i asked?

OpenStudy (anonymous):

sorry, so instead of answer in interval notations, i can just say x=-5 and im good